# Li Tana,b , Shengrong Wanga,c

### v1

@@ 5.2 Strong convergence rate
$\leq$ $C\int_{0}^{t}\sup_{0\leq u\leq s}\mathbb{E}|Y_{\Delta}^{i,N}(u)-X_{u}^{i,N}|^{p}{\rm d}s+C\Delta^{(1-\alpha)p}+C\Delta^{[(1-\alpha)\wedge H]p}.$
By Assumption 2.4 and Lemmas 5.1, 5.4, we get
$\mathbb{E}\bigg{(}\sup_{0\leq u\leq t}H_{3}(u)\bigg{)}$
$\leq$ $C\mathbb{E}\int_{0}^{t}|I(s)|^{p-2}[|Y_{\Delta}^{i,N}(s)-X_{s}^{i,N}|^{2}+|Y_{\Delta}^{i,N}(s-\tau)-X_{s-\tau}^{i,N}|^{2}+(\mathbb{W}_{q}(\mu^{Y,N}(s),\mu^{X,N}(s)))^{2}]{\rm d}s$
$+C\mathbb{E}\int_{0}^{t}|I(s)|^{p-2}|\theta b_{\Delta}(Y_{\Delta}^{i,N}(s),Y_{\Delta}^{i,N}(s-\tau),\mu^{Y,N}(s))\Delta||b(Y_{\Delta}^{i,N}(s),Y_{\Delta}^{i,N}(s-\tau),\mu^{Y,N}(s))$
$-b(X_{s}^{i,N},X_{s-\tau}^{i,N},\mu_{s}^{X,N})|{\rm d}s$
@@
$+C\Delta^{1-\alpha}\mathbb{E}\int_{0}^{t}|I(s)|^{p-2}[(1+|Y_{\Delta}^{i,N}(s)|^{l}+|Y_{\Delta}^{i,N}(s-\tau)|^{l}+|X_{s}^{i,N}|^{l}+|X_{s-\tau}^{i,N}|^{l})$
×$\times(|Y_{\Delta}^{i,N}(s)-X_{s}^{i,N}|+|Y_{\Delta}^{i,N}(s-\tau)-X_{s-\tau}^{i,N}|)+\mathbb{W}_{q}(\mu^{Y,N}(s),\mu_{s}^{X,N})]{\rm d}s$
$\leq$ $C\int_{0}^{t}\sup_{0\leq u\leq s}\mathbb{E}|Y_{\Delta}^{i,N}(u)-X_{u}^{i,N}|^{p}{\rm d}s+C\Delta^{(1-\alpha)(1+p)}.$
By the Young’s inequality, Assumption 2.5, we get
$\mathbb{E}\bigg{(}\sup_{0\leq u\leq t}H_{4}(u)\bigg{)}+\mathbb{E}\bigg{(}\sup_{0\leq u\leq t}H_{5}(u)\bigg{)}$
$\leq$ $C\mathbb{E}\int_{0}^{t}|I(s)|^{p}{\rm d}s+C\mathbb{E}\int_{0}^{t}(\mathbb{W}_{q}(\overline{\mu}^{Y,N}(s)),\mu^{Y,N}(s))^{p}{\rm d}s+C\mathbb{E}\int_{0}^{t}(\mathbb{W}_{q}(\mu^{Y,N}(s)),\mu_{s}^{X,N})^{p}{\rm d}s$
$\leq$ $C\int_{0}^{t}\sup_{0\leq u\leq s}\mathbb{E}|Y_{\Delta}^{i,N}(u)-X_{u}^{i,N}|^{p}{\rm d}s+C\Delta^{(1-\alpha)p}+C\Delta^{[(1-\alpha)\wedge H]p}.$
In the way in (3.3), by the Young’s inequality, we get
$\mathbb{E}\bigg{(}\sup_{0\leq u\leq t}|H_{6}(u)|^{p}\bigg{)}$
$\leq$ $C\mathbb{E}\int_{0}^{t}|I(s)|^{p}{\rm d}s+C\mathbb{E}\int_{0}^{t}(\mathbb{W}_{q}(\overline{\mu}^{Y,N}(s)),\mu^{Y,N}(s))^{p}{\rm d}s+C\mathbb{E}\int_{0}^{t}(\mathbb{W}_{q}(\mu^{Y,N}(s)),\mu_{s}^{X,N})^{p}{\rm d}s$
$\leq$ $C\int_{0}^{t}\sup_{0\leq u\leq s}\mathbb{E}|Y_{\Delta}^{i,N}(u)-X_{u}^{i,N}|^{p}{\rm d}s+C\Delta^{(1-\alpha)p}+C\Delta^{[(1-\alpha)\wedge H]p}.$
Combining $H_{1}-H_{6}$, we get
$\mathbb{E}\bigg{(}\sup_{0\leq u\leq t}|I(u)|^{p}\bigg{)}\leq C\int_{0}^{t}\mathbb{E}\bigg{(}\sup_{0\leq u\leq s}|Y_{\Delta}^{i,N}(u)-X_{u}^{i,N}|^{p}\bigg{)}{\rm d}s+C\Delta^{\alpha p}.$
Recalling the definition of $I(t)$ and by Assumption 2.3, we may compute
$|I(t)|^{p}\geq$ $2^{1-p}|Y_{\Delta}^{i,N}(t)-X_{t}^{i,N}-D(Y_{\Delta}^{i,N}(t-\tau))+D(X_{t-\tau}^{i,N})|^{p}$
@@
By Lemma 5.1, 5.4, we derive
$\mathbb{E}\bigg{(}\sup_{0\leq u\leq t}|X_{u}^{i,N}-Y_{\Delta}^{i,N}(u)|^{p}\bigg{)}\leq$ $\mathbb{E}\bigg{(}\sup_{0\leq u\leq t}|I(u)|^{p}\bigg{)}+C\Delta^{(1-\alpha)p}$
$\leq$ $C\int_{0}^{t}\mathbb{E}\bigg{(}\sup_{0\leq u\leq s}|Y_{\Delta}^{i,N}(u)-X_{u}^{i,N}|^{p}\bigg{)}{\rm d}s+C\Delta^{\alpha p}.$
Applying the Gronwall inequality, we can derive the assertion. ∎
@@
Let Assumptions 2.2-2.5 hold. For $p>1/H$ and $\overline{p}\geq 2(l+1)p$, it holds that
$\mathbb{E}\bigg{(}\sup_{t\in[0,T]}|X_{t}^{i}-Y_{t}^{i.N}|^{p}\bigg{)}\leq C\times\left\{\begin{array}[]{ll}N^{-\frac{1}{2}}+\Delta^{\alpha p},\ \ p>\frac{d}{2},\\ N^{-\frac{1}{2}}\log(1+N)+\Delta^{\alpha p},\ p=\frac{d}{2},\\ N^{-\frac{p}{d}}+\Delta^{\alpha p},\ p\in[2,\frac{d}{2}),\end{array}\right.$

### v2

@@ 5.2 Strong Convergence Rate
$\leq$ $C\int_{0}^{t}\mathbb{E}\left(\sup_{0\leq u\leq s}|Y_{\Delta}^{i,N}(u)-X_{u}^{i,N}|^{p}\right){\rm d}s+C\Delta^{(1-\alpha)p}+C\Delta^{[(1-\alpha)\wedge H]p}.$
Again, by Assumption 2.4 and Lemmas 5.1, 5.4
$\mathbb{E}\bigg(\sup_{0\leq u\leq t}Z_{3}(u)\bigg)$
$\leq$ $C\mathbb{E}\int_{0}^{t}|\Lambda(s)|^{p-2}[|Y_{\Delta}^{i,N}(s)-X_{s}^{i,N}|^{2}+|Y_{\Delta}^{i,N}(s-\tau)-X_{s-\tau}^{i,N}|^{2}+\mathbb{W}_{2}^{2}(\mu^{Y,N}(s),\mu^{X,N}(s))]{\rm d}s$
$+C\mathbb{E}\int_{0}^{t}|\Lambda(s)|^{p-2}|\theta b_{\Delta}(Y_{\Delta}^{i,N}(s),Y_{\Delta}^{i,N}(s-\tau),\mu^{Y,N}(s))\Delta|$
$\cdot|b(Y_{\Delta}^{i,N}(s),Y_{\Delta}^{i,N}(s-\tau),\mu^{Y,N}(s))-b(X_{s}^{i,N},X_{s-\tau}^{i,N},\mu_{s}^{X,N})|{\rm d}s$
@@
$+C\Delta^{1-\alpha}\mathbb{E}\int_{0}^{t}|\Lambda(s)|^{p-2}[(1+|Y_{\Delta}^{i,N}(s)|^{l}+|Y_{\Delta}^{i,N}(s-\tau)|^{l}+|X_{s}^{i,N}|^{l}+|X_{s-\tau}^{i,N}|^{l})$
$\times(|Y_{\Delta}^{i,N}(s)-X_{s}^{i,N}|+|Y_{\Delta}^{i,N}(s-\tau)-X_{s-\tau}^{i,N}|)+\mathbb{W}_{2}(\mu^{Y,N}(s),\mu_{s}^{X,N})]{\rm d}s$
$\leq$ $C\int_{0}^{t}\mathbb{E}\left(\sup_{0\leq u\leq s}|Y_{\Delta}^{i,N}(u)-X_{u}^{i,N}|^{p}\right){\rm d}s+C\Delta^{(1-\alpha)(1+p)}.$
Further, the Young’s inequality, Assumption 2.5, and Lemmas 5.1, 5.5 lead to
$\mathbb{E}\bigg(\sup_{0\leq u\leq t}Z_{4}(u)\bigg)+\mathbb{E}\bigg(\sup_{0\leq u\leq t}Z_{5}(u)\bigg)$
$\leq$ $C\mathbb{E}\int_{0}^{t}|\Lambda(s)|^{p}{\rm d}s+C\mathbb{E}\int_{0}^{t}(t-s)^{2H-1}\mathbb{W}_{2}^{p}(\overline{\mu}^{Y,N}(s),\mu^{Y,N}(s)){\rm d}s$
$+C\mathbb{E}\int_{0}^{t}(t-s)^{2H-1}\mathbb{W}_{2}^{p}(\mu^{Y,N}(s),\mu_{s}^{X,N}){\rm d}s$
$\leq$ $C\int_{0}^{t}\mathbb{E}\left(\sup_{0\leq u\leq s}|Y_{\Delta}^{i,N}(u)-X_{u}^{i,N}|^{p}\right){\rm d}s+C\int_{0}^{t}(t-s)^{2H-1}\mathbb{E}\bigg(\sup_{0\leq u\leq s}|Y_{\Delta}^{i,N}(u)-X_{u}^{i,N}|^{p}\bigg){\rm d}s$
$+C\Delta^{(1-\alpha)p}+C\Delta^{[(1-\alpha)\wedge H]p}.$
In the similar way as achieving (3.6), the Young’s inequality results in
$\mathbb{E}\bigg(\sup_{0\leq u\leq t}|Z_{6}(u)|^{p}\bigg)\leq C\Delta^{(1-\alpha)p}+C\Delta^{[(1-\alpha)\wedge H]p}$
$+C\epsilon\mathbb{E}\bigg(\sup_{0\leq u\leq t}|\Lambda(u)|^{p}\bigg)+C\int_{0}^{t}\mathbb{E}\left(\sup_{0\leq u\leq s}|Y_{\Delta}^{i,N}(u)-X_{u}^{i,N}|^{p}\right){\rm d}s.$
Collecting the above estimates $Z_{1}-Z_{6}$ together
$\mathbb{E}\bigg(\sup_{0\leq u\leq t}|\Lambda(u)|^{p}\bigg)\leq$ $C\int_{0}^{t}\mathbb{E}\bigg(\sup_{0\leq u\leq s}|Y_{\Delta}^{i,N}(u)-X_{u}^{i,N}|^{p}\bigg){\rm d}s$
$+C\int_{0}^{t}(t-s)^{2H-1}\mathbb{E}\left(\sup_{0\leq u\leq s}|Y_{\Delta}^{i,N}(u)-X_{u}^{i,N}|^{p}\right){\rm d}s+C\Delta^{\alpha p}.$
Recalling the definition of $\Lambda(t)$ and using Assumption 2.3, we may compute
$|\Lambda(t)|^{p}\geq$ $2^{1-p}|Y_{\Delta}^{i,N}(t)-X_{t}^{i,N}-D(Y_{\Delta}^{i,N}(t-\tau))+D(X_{t-\tau}^{i,N})|^{p}$
@@
This gives from Lemmas 5.1, 5.4 that
$\mathbb{E}\bigg(\sup_{0\leq u\leq t}|X_{u}^{i,N}-Y_{\Delta}^{i,N}(u)|^{p}\bigg)\leq\mathbb{E}\bigg(\sup_{0\leq u\leq t}|\Lambda(u)|^{p}\bigg)+C\Delta^{(1-\alpha)p}$
$\leq$ $C\Delta^{\alpha p}+C\int_{0}^{t}\mathbb{E}\bigg(\sup_{0\leq u\leq s}|Y_{\Delta}^{i,N}(u)-X_{u}^{i,N}|^{p}\bigg){\rm d}s$
$+C\int_{0}^{t}(t-s)^{2H-1}\mathbb{E}\left(\sup_{0\leq u\leq s}|Y_{\Delta}^{i,N}(u)-X_{u}^{i,N}|^{p}\right){\rm d}s.$
Since the function $E_{2H,1}((\Gamma(2H))^{1/2H}t)$ converges, we can utilize Lemma 2.8 to get the assertion. ∎
@@
Let Assumptions 2.2-2.5 hold. Fix any $\overline{p}>4(l+1)$. Then, for $p\in[2,\frac{\overline{p}}{2(l+1)})$, it holds that
$\mathbb{E}\bigg(\sup_{t\in[0,T]}|X_{t}^{i}-Y_{\Delta}^{i,N}(t)|^{p}\bigg)\leq C\times\left\{\begin{array}[]{ll}N^{-\frac{1}{2}}+\Delta^{\alpha p},\ \ p>\frac{d}{2},\\ N^{-\frac{1}{2}}\log(1+N)+\Delta^{\alpha p},\ p=\frac{d}{2},\\ N^{-\frac{p}{d}}+\Delta^{\alpha p},\ p\in[2,\frac{d}{2}),\end{array}\right$.
@@
Remark 5.8.
This paper introduces a tamed theta EM scheme for a class of McKean-Vlasov SDEs driven by fractional Brownian motions. In contrast to the existing literature, the proposed framework generalizes several tamed EM schemes studied in works such as [13] and [37]. The convergence rate is rigorously derived by employing fractional Itô calculus, the fractional BDG inequality, and other pertinent inequalities. This methodology provides a new perspective for handling this category of fractional McKean-Vlasov equations.